(x-2)^(2/3)+(x-2)^(1/3)-12=0
the fractions are exponents

Found 2 solutions by lynnlo, Alan3354:
Answer by lynnlo(4176)

is there a number that goes after ^ sign

Answer by Alan3354(69443)

(x-2)^(2/3)+(x-2)^(1/3)-12=0
=============
Sub y for (x-2)^(1/3)
---
y^2 + y - 12 = 0
(y-3)*(y+4) = 0
y = 3
x^(1/3) = 3
x = 27
==============
y = -4
x^(1/3) = -4
x = -64